OpenStudy (anonymous):

Place the symbols to enter the trigonometric equation (sin, tan, or cos) you would use to find x in the following right triangles. Do not solve the equations. i will draw

OpenStudy (anonymous):



OpenStudy (anonymous):

sin20=o/h , cos20=a/h tan 20=o/a

OpenStudy (anonymous):

which sides are you given and apply that trig function

OpenStudy (anonymous):

how do you know what o/h a/h and o/a mean

OpenStudy (anonymous):

SOH CAH TOA is the method that you should memorized. SOH stands for sin(...)=Opposite/Hypotenuse CAH stands for cos(...)=Adjacent/Hypotenuse TOA stands for tan(...)=Opposite/Adjacent Make sense? If you're ever going to be taking any more math classes in your life you really need to memorize that -- super helpful! In your specific problem you have x and 12 as your sides and no hypotenuse. So which trig function can you use?

OpenStudy (anonymous):

*memorize

OpenStudy (anonymous):

thanks